Supplemental Operations Report – April 27th

Supplemental Operations Report 27 April 2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: HabCar rear window fell out. The hinges broke and no one reported it.

Action taken for non-nominal systems: I removed the window completely. It will need new hinges to remount the window.

Generator: Still being manually run at night

ScienceDome Dual Split: Off

Solar—Charge controllers off at night

Solar— Nominal

Notes on power system: Nothing to report

Propane Reading, station tank – 75 %

Propane Reading, director tank— 82 %

Propane Reading, intern tank— 70 %

Propane Reading, generator—82 %

Ethanol Free Gasoline – 0 gallons.

Water (Outpost tank) – 500 gallons

Sojourner rover used: yes

Hours: 168.8

Beginning charge: 100

Ending charge: 88, now 100

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 350.1, 350.2, 300): none

Reason for use: n/a

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0 hours

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town for garbage and water. I was extremely disappointed to discover that the rear window hinges had broken and no one told me. I saw that it was broken but had no idea as to when so I drove with it broken and the window fell out. It did not break as it was being held in place by one of the side rods. I was able to get it inside and drove back to the station and popped the other rod out with a crowbar. I will need to order new hinges and then the window can be remounted, but until then the back window will be missing. I stored it in the RAM.

CrewCar used and why, where? The noise I was hearing from the front of the car continues as well as the occasional loss of power steering. I ordered power steering fluid, even though when I had the car serviced they reported all fluids nominal. I’ll see if that helps, if not we will need to take it in to see what is wrong either during the Work Party this spring or next fall.

Luna used and why, where? Yes, to town, but now parked until I can look at the tire pressure.

Campus-wide inspection, if action taken, what and why? Nothing to report

General notes and comments: Nothing to report

Summary of general operations: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of Outpost operations: I opened the box with the new pump to find it was a resold return. Since you can use a submersible pump in all kinds of water, I did not open it up but rather opted to return it. I can’t risk contamination of our drinking water again. This means that I still don’t have a way to get water out of the storage tank and so made an emergency trip to town for water I could pump into my trailer from the HabCar tank. I ordered another pump.

I removed the fence on the east side of the yard so that the propane truck can make it. Propane was delivered today and now that I know how much space he will need, I will rebuild the fence. This will also allow us to park a vehicle closer to the trailer, which will be helpful at times.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Operations Report – April 27th

Crew 265 Operations Report Sol 3 27-04-2022

SOL:3

Name of person filing report: Dave Laude/Sergii Iakymov

Non-nominal systems: Furnace, Suit #8. Percy’s left rear tail light is non-functional, but the brake light is working.

Notes on non-nominal systems: Furnace is functional and we remain cautious over limiting to early morning only. Infrequently used now. Suit #8 has blown a fuse during our mission while on the table.

ROVERS
Spirit rover used: Yes

Hours: 174.3
Beginning charge: 100
Ending charge: 100!
Currently charging: Yes

Opportunity rover used: No
Hours:
Beginning charge:
Ending charge:
Currently charging:

Curiosity rover used: No
Hours:
Beginning charge:
Ending charge:
Currently charging:

Perseverance rover used: Yes
Hours: 236.5
Beginning charge: 100
Ending charge: 89
Currently charging: Yes

General notes and comments:

Summary of Hab operations:

WATER USE:
Water (static tank): Not full. ~410 gallons.
Water (loft tank): 30 gallons
Water Meter: 158117.6 units

Static to Loft Pump used – Yes
Static tank pipe heater (on or off): Off
Static tank heater (On or off): Off
Toilet tank emptied: Yes

Summary of internet: Nominal

Summary of suits and radios: Batteries charging.
Suite #8 has been repaired (blown fuse has been replaced; loose wires soldered and isolated with electrical tape ). Additionally, we discovered a loose charging port in the same suite. After some adjustments and repairs the suit is functional again. The receiving connectors on the suit were slightly expanded open by repetitive insertions and removals of the charger plug making the charging dependent upon adjusting the plug to just the correct position for proper contact. Now just plug in and charge without fussing with it, for now. There are two other one-piece suits that require attention that we plan to get at soon.

Summary of GreenHab operations: NA

WATER USE: gallons

Heater: On or Off

Supplemental light: On or Off

Harvest:

Summary of ScienceDome operations: NA

Dual split: Heat or AC, On or Off

Summary of RAM operations: Used soldering iron, solder and tape to repair suit #8.

Summary of any observatory issues: NA

Summary of health and safety issues: Per HSO report one person is prohibited from EVA due to high blood pressure and another is being monitored for bp and heart rate a few times a day.

Questions, concerns and requests to Mission Support: (note trash pickup)

Getting hot in hab upper deck. What is the status of the swamp cooler water supply and any idea how much water it could consume?

Operations Report – April 26th

Crew 265 Operations Report Sol 2 26-04-2022

SOL:2

Name of person filing report: Dave Laude/Sergii Iakymov

Non-nominal systems: Furnace, suit #8

Notes on non-nominal systems: Furnace is functional and we remain cautious over its limiting to early morning only. The suit has blown a fuse during our mission while on the table.

ROVERS

Spirit rover used: Yes

Hours: 173.6

Beginning charge: 100%

Ending charge: 86%

Currently charging: Yes

Opportunity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Perseverance rover used: Yes

Hours: 235.7

Beginning charge: 100

Ending charge: 68

Currently charging: Yes

General notes and comments:
Summary of Hab operations:
WATER USE:

Water (static tank): Not full. ~445 gallons.

Water (loft tank): 25 gallons

Water Meter: 158080.8units
Static to Loft Pump used – No

Static tank pipe heater (on or off): Off

Static tank heater (On or off): Off

Toilet tank emptied: No
Summary of internet: Nominal

Summary of suits and radios: Batteries charging.

Suite #8 has a blown fuse and loose wires. One of the suit chargers has a loose wire as well. Both will be repaired, but the source of the blown fuse is a mystery at the moment. We have a bag of spares.

Two suit batteries labeled as bad were tested and had low capacity at 2.5-3.5 A-hr. Two unlabeled and non-boxed batteries were tested at 2.9-4.7 A-hr. There are more unlabeled that I assume are deficient and not planning to test, but can test them if needed. I am now testing a boxed and apparently new battery.

Operations Report – April 25th

Crew 265 Operations Report Sol 1 25-04-2022

SOL: 1

Name of person filing report: Dave Laude/Sergii Iakymov

Non-nominal systems: Furnace. Curiosity Rover.

Notes on non-nominal systems: Furnace this morning worked properly. The stuck electrical contact seems to have come unstuck on the pressure sensor for now.

Curiosity rover ended with 90% charge after a short drive while others for the same drive were 96-98%. Started at 100%. Driver insisted the data is correct.

ROVERS

Spirit rover used: Yes

Hours: 173.1

Beginning charge: 100

Ending charge: 98

Currently charging: Yes

Opportunity rover used: Yes

Hours: Not recorded

Beginning charge: 100

Ending charge: 96

Currently charging: Yes

Curiosity rover used: Yes

Hours: 186.0

Beginning charge: 100

Ending charge: 90

Currently charging: Yes

Perseverance rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging:

General notes and comments:

Summary of Hab operations: Trained crew on EVA suits, radios and rovers then sim began at noon. Some food supplies were absent. Requested more.

WATER USE:

Water (static tank): Not full. ~445 gallons.

Water (loft tank): 55 gallons

Water Meter: 01580472 units

Static to Loft Pump used – Yes

Static tank pipe heater (on or off): Off

Static tank heater (On or off): Off

Toilet tank emptied: Yes

Summary of internet: Nominal

Summary of suits and radios: Batteries charging. Test of EVA suits (1-6, 8, 9) battery capacity by running fans for 3 hour 40 minutes has shown all to be in good condition by dropping in open circuit voltage to 12.58-12.93V. The remaining two one-piece suits were not on charge when we arrived and have battery voltages ~3V. Will investigate soon.

Summary of GreenHab operations: Not used, nothing to report

WATER USE: N/A

Heater: OFF

Supplemental light: OFF

Harvest: N/A

Summary of ScienceDome operations: Not used, nothing to report

Dual split: Heat or AC, Off

Summary of RAM operations: Tools used for furnace work. DMMs used for battery checks.

Summary of any observatory issues: Won’t be using, nothing to report

Summary of health and safety issues: See HSO Report

Questions, concerns and requests to Mission Support: one bag of burnable is ready for pick up in the engineering airlock.

Water (static tank): Not full. ~435 gallons.

Crew 265 Operations Report 24 APR 2022

[Crew 265 Operations Report April 24th]

Crew 265 Operations Report 24-04-2022

SOL: 0

Name of person filing report: Dave Laude

Non-nominal systems: Heating system.

Notes on non-nominal systems: The heating system is now functional after non-functioning with previous crew. I used the first new thermostat that was ordered as that is correct for this system. However, it does not have fan switches like the original did. The main issue is that the exhaust gas fan to the outside cycles on and off because the fan operation pressure sensing switch is stuck closed and so the furnace senses this as a fault. For operation I have to fake the opening and closing during start-up as it is being checked by the electronics. We only use the furnace in the morning for a few hours if very cold in the hab and when I can tend to it. Warm temps are coming soon and so we may not need it. It will not be run overnight as is. No replacement part is planned to be ordered because the entire furnace is to be replaced after this crew season is over.

ROVERS

Spirit rover used: No

Hours: (before EVA)

Beginning charge: (Before EVA)

Ending charge: (On return from EVA, before recharging)

Currently charging: Yes

Opportunity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Perseverance rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

General notes and comments:

Summary of Hab operations:

WATER USE:

Water (static tank): Not full. ~480 gallons. Does this warrant a special trip for water

Water (loft tank): 20 gallons

Water Meter: 0158026.1 units

Static to Loft Pump used – No

Static tank pipe heater (on or off): Off

Static tank heater (On or off): Off

Toilet tank emptied: No

Summary of internet: Awesome!

Summary of suits and radios: Batteries charging. One radio in RAM seems to have a very weak battery battery. Will need to test it. Radio replaced with one from lower hab deck. Testing of some currently unused EVA suit battery capacity in inventory has started. Scott Davis has asked for this on a sample basis to confirm or not that batteries labeled dead are of no use. Will start some radio battery capacity testing soon.

Summary of GreenHab operations: Not used, nothing to report

WATER USE: gallons

Heater: On or Off

Supplemental light: On or Off

Harvest:

Summary of ScienceDome operations: Not used, nothing to report

Dual split: Heat or AC, Off

Summary of RAM operations: Tools used for furnace work. DMMs used for battery checks.

Summary of any observatory issues: Won’t be using, nothing to report

Summary of health and safety issues: Star Link cable has been taped to the floor on the lower deck to prevent tripping upon it.

Questions, concerns and requests to Mission Support: (note trash pickup)

Water (static tank): Not full. ~85% or ~470 gallons. Maybe needs another 6-8". Does this warrant a special trip for water tomorrow or can someone deliver some during the mission?

Minutes ago we smelled natural gas or propane on the lower deck only. The gas detector alarm was not set off. Manual detectors read null and the gas dissipated in about 10 minutes. This same situation has occurred on other crews. We will check periodically for the same occurance.

Supplemental Operations Report April 24th

Supplemental Operations Report 24 April 2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: n/a

Generator: Still being manually run at night from about 9 pm to 7:30 am

ScienceDome Dual Split: Off

Solar—Charge controllers off at night.

Solar— Nominal

Notes on power system: Nothing to report

Propane Reading, station tank – 57 %

Propane Reading, director tank— 58 %

Propane Reading, intern tank— 68 %

Propane Reading, generator—43 %

Ethanol Free Gasoline – 0 gallons.

Water (Outpost tank) – 500 gallons

Sojourner rover used: no

Hours: 167.7

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: All rovers were serviced. The batteries were cleaned, neutralized and the cells of the batteries were refilled.

Spirit showed some corrosion and batteries were evenly depleted of water.

Curiosity showed some corrosion and batteries were unevenly depleted of water. The driver’s side had water levels equal to that of Spirit, but the passenger’s side were much more depleted. I think they were accidentally not filled the last time.

Percy showed limited corrosion and batteries were evenly depleted of water. One battery took some effort to refill due to a rusted bolt.

Opportunity showed some corrosion and batteries were evenly depleted of water.

Sojourner showed heavier corrosion and batteries were evenly depleted of water.

It took about 70 gallons of water to clean them and 13 gallons of distilled water to fill them. They will need to be serviced again during the work party.

ATV’s Used: (Honda, 350.1, 350.2, 300): none

Reason for use: n/a

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0 hours

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town for garbage and water

CrewCar used and why, where? Yes, to Grand Junction. I fixed the broken piece on the interior driver’s side by the windshield. It just needed to pushed securely back into place.

Luna used and why, where? Yes, to Denver. Got a warning that the tire sensor is inoperable and needs to be reset. I will work on figuring this one out. Also the tires showed four different pressures right before this, so I will adjust them as needed.

Campus wide inspection, if action taken, what and why? Nothing to report

General notes and comments: Dave and Sergii got here on Saturday and helped me do things before the arrival of the crew. This was deeply appreciated.

Summary of general operations: Nothing to report

Summary of internet: We removed the multiple sets of HudgesNet and old internet wiring throughout the Hab. We also removed the blue ethernet wire that had been put in place for the webcams that could not be used. Things look much neater as a result.

Summary of suits and radios: Dave is working with Scott on some troubleshooting of some of the suits and batteries.

Summary of Hab operations: Dave did some troubleshooting of the furnace and it is semi-operational. It can be turned on to warm the Hab but it is not completely safe as he had to bypass an exhaust sensor to do so. In addition, the new thermostat does not have an on/off switch and turns on at 50 degrees, so it will need to be disabled after each use by disconnecting the wires. This is a good enough work around for this crew, and as they are the last crew that will need it before a new furnace is installed, it will work. They will not run the furnace at night and in a few days it should be warm enough to retire the furnace forever. Dave also reset the water alarm in the loft tank.

Summary of Outpost operations: The new pump and hose have arrived for the water tank and I will set it up when I need more water.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: I dropped a 10mm wrench from the newer set between the batteries of Percy and could not retrieve it. There is an older 10 mm single wrench that can replace it until we can get it out from where it is wedged.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

[end]

Operations Report – April 20th

Crew 245 Operations Report 20-04-2022
SOL: 10
Name of the person filing report: Luca Rossettini

Non-nominal systems: Heating system.
Notes on non-nominal systems: –

ROVERS

Spirit rover used: Yes
Hours: 172.4
Beginning charge: 100%
Ending charge: 80%
Currently charging: Yes

Opportunity rover used: Yes
Hours: 95.6
Beginning charge: 100%
Ending charge: 95%
Currently charging: Yes

Curiosity rover used: Yes
Hours: 184.9
Beginning charge: 100%
Ending charge: 78%
Currently charging: Yes

Perseverance rover used: No
Hours: 235.6
Beginning charge: 100%
Ending charge: -%
Currently charging: –

General notes and comments: None

Summary of Hab operations:

WATER USE
Water (static tank): 342 gallons
Water (loft tank): 12 gallons
Water Meter: 157861.9
Static to Loft Pump used: Yes
Static tank pipe heater: Off
Static tank heater: Off
Toilet tank emptied: No

Summary of internet: nothing to report

Summary of suits and radios: All suits nominal. We found a charger plug not properly connected and fixed the problem. Because of that, the batteries that we thought to be exhausted at the beginning of the simulation, in fact, were not. We put them in charge and they are working properly. The LEDs on the chargers stay on even if the plug is not properly connected to the wall and the suit battery is connected to the charger. This behavior could be misleading.

Summary of GreenHab operations:

WATER USE: 0 gallons
Heater: Off
Supplemental light: Off
Harvest:

Summary of ScienceDome operations: Nothing to report

Dual split (Heat/AC): Off

Summary of RAM operations: nothing to report

Summary of any observatory issues: Not used, nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: The LEDs on the chargers stay on even if the plug is not properly connected to the wall and the suit battery is connected to the charger. This behavior could be misleading.

Operations Report – April 19th

Crew 245 Operations Report 19-04-2022
SOL: 9
Name of the person filing report: Luca Rossettini

Non-nominal systems:
Suits 3 and 1.
Heating system.

Notes on non-nominal systems: we will proceed implementing the instructions provided by MDRS Support.

ROVERS
Spirit rover used: Yes
Hours: 171.8
Beginning charge: 100%
Ending charge: 81%
Currently charging: No
Opportunity rover used: No
Hours: 95.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes
Curiosity rover used: No
Hours: 184.4
Beginning charge: 100%
Ending charge: 98%
Currently charging: Yes
Perseverance rover used: No
Hours: 235.6
Beginning charge: 100%
Ending charge: –
Currently charging: –
General notes and comments: All the rovers were serviced and back at the hub, with the exception of Perseverance, still under maintenance, and without and estimated time of completion.
Summary of Hab operations:
WATER USE
Water (static tank): 351 gallons
Water (loft tank): 12 gallons
Water Meter: 157819.9
Static to Loft Pump used: Yes
Static tank pipe heater: Off
Static tank heater: Off
Toilet tank emptied: No
Summary of internet: nothing to report
Summary of suits and radios: Suit No. 3 is not yet working. We are still investigating the potential issue following the instructions received.
Suit No. 1 is also not fully working, and is showing two issues: a. the air is flowing only from one side of the backpack; b. rotating the air intensity regulator, it is not fully possible to regulate the flow: either the intensity is very low for most of the regulations, or we have to turn it to max. Even in these cases we will follow the instructions given for Suit 3, unless stopped by MDRS Support.
Summary of GreenHab operations:
WATER USE: 0 gallons
Heater: Off
Supplemental light: Off
Harvest:
Summary of ScienceDome operations: Nothing to report
Dual split (Heat/AC): Off
Summary of RAM operations: nothing to report
Summary of any observatory issues: Not used, nothing to report
Summary of health and safety issues: nothing to report
Questions, concerns and requests to Mission Support: –

Supplemental Operations Report – April 17th

Supplemental Operations Report 17 April 2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Water pump in outpost storage tank, Hab furnace

Action taken for non-nominal systems: Ordered new pump and hose to temporarily use for the remainder of this season and we can install a new permanent pump in the fall. I ordered the wrong thermostat for the furnace but am going to just retire it as we will be getting a new one in the fall and we just experienced probably the final cold weather of the season.

Generator: Still being manually run at night.

ScienceDome Dual Split: Off

Solar—Charge controllers off at night.
Solar— Nominal

Notes on power system: Nothing to report

Propane Reading, station tank – 58 %
Propane Reading, director tank— 50%
Propane Reading, intern tank— 69 %
Propane Reading, generator— 68 % (I think—I forgot between there and here!)

Ethanol Free Gasoline – 0 gallons.

Water (Outpost tank) – 350 gallons

Sojourner rover used: no
Hours: 167.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Notes on rovers: Oppy is starting to overheat as it gets hot. I have moved it off campus for service and once I have done it, I will move it back for short and/or slow trips.

I’m all set up to service the rovers tomorrow. The batteries will be cleaned and neutralized, then the cells refilled with distilled water.

ATV’s Used: (Honda, 350.1, 350.2, 300): none
Reason for use: n/a
Oil Added? No.
ATV Fuel Used: 0 gallons
# Hours the ATVs were Used: 0 hoursNotes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town for garbage and water

CrewCar used and why, where? Yes, to town for packages

Luna used and why, where? Yes, to town. I changed the cabin and regular air filters.

Campus-wide inspection, if action is taken, what and why? Things are holding up very well in the wind this year. We haven’t had many windstorms this season though, so that’s probably part of it. The Hab has gotten the worse of it and some of the outer walls have damage.

General notes and comments: There will not be a Supplemental Operations Report this Wednesday.

Summary of general operations: Nothing to report
Summary of internet: Nothing to report
Summary of suits and radios: Nothing to report
Summary of Hab operations: Nothing to report

Summary of Outpost operations: New toilet installed and working well. But then the pump in the now clean water storage tank stopped working. I moved the plants from the intern trailer to the ScienceDome.

Summary of GreenHab operations: The windstorms have deposited a lot of sand in the GreenHab.
Summary of ScienceDome operations: Nothing to report
Summary of RAM operations: Nothing to report
Summary of any observatory issues: Nothing to report
Summary of health and safety issues: Nothing to report